# Axial Neural Networks for Dimension-Free Foundation Models

### v1

@@ Appendix B Proofs of Theorems
Using this identity and Eq.˜45, permutation of the input yields
$\sum_{\pi^{\prime}(l)=1}^{K}\text{T}_{(K)(\pi^{\prime}(l))}\Pi^{\prime}\boldsymbol{h}^{\prime}_{d_{l}}$ $=\sum_{\pi^{\prime}(l)=1}^{K}\Pi^{\prime}\text{T}_{(K)(l)}\boldsymbol{h}^{\prime}_{d_{l}}=\sum_{\pi^{\prime}(l)=1}^{K}\text{T}_{(K)(\pi(l))}\Pi\boldsymbol{h}^{\prime}_{d_{l}}$ (48)
$=\sum_{\pi^{\prime}(l)=1}^{K}\Pi\text{T}_{(K)(l)}\boldsymbol{h}^{\prime}_{d_{l}}=\Pi\sum_{l=1}^{K}\text{T}_{(K)(l)}\boldsymbol{h}^{\prime}_{d_{l}}=\Pi\boldsymbol{h}^{\prime\prime},$
@@ Appendix C X-CViT and X-MPP
$\boldsymbol{h}_{d_{i}}=\frac{1}{K}\sum_{j\neq i}^{K}\text{T}_{(K)(i)}\left(\text{T}_{(K-1)(j)}\left(\operatorname{MaxPool}_{1,\ldots,K-2}\left(\operatorname{Conv2D}_{K-1,K}\left(\text{T}_{(j)(K-1)}\text{T}_{(i)(K)}(\boldsymbol{x})\right)\right)\right)\right),$ (50)
where $\text{T}_{(m)(n)}(\boldsymbol{x})$ denotes the transpose between the $m$-th and $n$-th axes, $\text{T}_{(j)(K-1)}\text{T}_{(i)(K)}$ is used to apply $\operatorname{Conv2D}_{K-1,K}$ along the $i$-th and $j$-th axes, and $\text{T}_{(K)(i)}\text{T}_{(K-1)(j)}$ reorders the axes back to enable aligned feature aggregation through the summation.

### v2

@@ B.4 Axes-Permutation Equivariance of Pooling Layer
Using this identity and Eq.˜40, permutation of the input yields
$\sum_{\pi^{\prime}(l)=1}^{K}\text{T}_{(\pi^{\prime}(l))(K)}^{-1}\Pi^{\prime}\boldsymbol{h}^{\prime}_{d_{l}}$ $=\sum_{\pi^{\prime}(l)=1}^{K}\Pi^{\prime}\text{T}_{(l)(K)}^{-1}\boldsymbol{h}^{\prime}_{d_{l}}=\sum_{\pi^{\prime}(l)=1}^{K}\text{T}_{(\pi(l))(K)}^{-1}\Pi\boldsymbol{h}^{\prime}_{d_{l}}$ (43)
$=\sum_{\pi^{\prime}(l)=1}^{K}\Pi\text{T}_{(l)(K)}^{-1}\boldsymbol{h}^{\prime}_{d_{l}}=\Pi\sum_{l=1}^{K}\text{T}_{(l)(K)}^{-1}\boldsymbol{h}^{\prime}_{d_{l}}=\Pi\boldsymbol{h}^{\prime\prime},$
@@ Appendix C X-CViT and X-MPP
$\boldsymbol{h}_{d_{i}}=\frac{1}{K}\sum_{j\neq i}^{K}\left(\text{T}_{(K-1)(j)(K)(i)}\left(\operatorname{MaxPool}_{1,\ldots,K-2}\left(\operatorname{Conv2D}_{K-1,K}\left(\text{T}_{(i)(K)(j)(K-1)}(\boldsymbol{x})\right)\right)\right)\right),$ (45)
where $\text{T}_{(i)(K)(j)(K-1)}$ is used to apply $\operatorname{Conv2D}_{K-1,K}$ along the $i$-th and $j$-th axes, and $\text{T}_{(K-1)(j)(K)(i)}$ reorders the axes back to enable aligned feature aggregation through the summation.